OpenStudy (anonymous):

the length of the top of the table is 4 m greater than the width. the area is 60 m^2. find the dimensions of the table.

OpenStudy (anonymous):

w = l - 4 l(l-4) = 60 l^2 - 4l - 60 = 0 (l +6)(l - 10) = 0 l = 10 w = 6 length = 10m width = 6 m

OpenStudy (anonymous):

thank you very much

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

plz click on good answer to close the question